OpenStudy (anonymous):

What is the cost of $50 jeans and a 7% sales tax

OpenStudy (anonymous):

You need to find 7% of 50, and then add it on to the original 50 dollar cost. \[\ 50\times \frac{ 7 }{ 100 }\] \[\frac{ 350 }{ 100 }\] So 7% of $50 is $3.50. If you add that up, you get that the final cost is $53.50.